OpenStudy (anonymous):

If A + B = 4/pi, then value of (1 + tan A) (1 + tan B) equals

OpenStudy (anonymous):

A+B=pi/4 tan(A+B)=tan pi/4 [Multiplying by tan on both sides] (tan A + tan B)/( 1 - tan A tan B) = 1 tan A + tan B = 1 - tan A tan B 1 + tan A + tan B + tan A tan B = 2 (1+tan A) (1+tan B) = 2